# Stable cohomology of universal character varieties

### v1

@@ 4.1. Monodromy representations
The universal $\operatorname{PGL}_{n}$ bundle on $S_{g}\times M_{B}^{\operatorname{GL}}(S_{g,1},n,d)$ described in [HT04, §1] is the pullback of $\mathcal{P}$ by the quotient map $M_{B}^{\operatorname{GL}}(S_{g,1},n,d)\to\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$, as is the universal $\operatorname{PGL}_{n}$-bundle on $S_{g}\times M_{B}^{\operatorname{SL}}(S_{g,1},n,d)$ in [HRV08, §4.1].
@@
The generation statement is a consequence of Markman’s tautological generation result for $M_{\mathrm{Dol}}^{\operatorname{GL}}(C,n,d)$ [Mar02, Theorem 7], as reinterpreted on the Betti side by Hausel and Thaddeus [HT04, (5.1)]. Picking a homogeneous basis of $H^{*}(S_{g};\mathbb{Q})$, there are no relations in degrees $\leq 2(g-1)(n-1)+2$ among the images of these elements in $V_{*,j}$ by [HRV08, Lemma 4.1.12]. ∎
@@
In later work [Mar07, Theorem 3], Markman strengthened the generation result from rational to integral cohomology. In particular, in degrees $\leq 2(g-1)(n-1)+2$ one concludes that $H^{*}(M^{\operatorname{GL}}_{B}(S_{g,1},n,d);\mathbb{Z})$ is torsion-free. Then using the quotient description of Proposition 2.5, we conclude that the cohomology groups of the fibers $H^{*}(M^{\operatorname{SL}}_{B}(S_{g,1},n,d);\mathbb{Z}[n^{-1}])$ and $H^{*}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Z}[n^{-1}])$ are torsion-free in the same degree range. However, in order to deduce a result about integral cohomology of the universal spaces, one would also need torsion-freeness results for the various cohomology groups of $\mathcal{M}_{g}$ and $\mathcal{M}_{g,1}$ with symplectic coefficients.
@@
We follow the strategy of [HT04, (5.1)]. For any element of $\operatorname{Mod}_{g}$, we may choose a representative diffeomorphism $f$ of $S_{g}$ fixing a chosen point $p$. It is then straightforward to check that $\mathcal{P}$ and $f^{*}\mathcal{P}$ are both $\operatorname{PGL}_{n}$-equivariant flat $\operatorname{PGL}_{n}$-bundles over $S_{g}\times X_{g,n,d}$ trivialized over $\{p\}\times X_{g,n,d}$ such that the monodromy representation of the restriction to $S_{g}\times\{\phi\}$ is given by $\phi$, and thus $\mathcal{P}\cong f^{*}\mathcal{P}$.
@@
are canonical, and in particular invariant under homeomorphisms of $S_{g}$. (See [She17] for more discussion.)
@@ 4.2. Stable rational cohomology of $\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)$
with $s_{i}=(1+2i)r_{i}^{\prime}$ for some $r_{i}^{\prime}\geq r_{i}$ satisfying $r_{i}^{\prime}\equiv r_{i}$ mod $2$. Such a term has highest weight $\sum\langle 1^{r_{i}}\rangle$, and so each irreducible piece $S_{\langle\lambda\rangle}(V_{g})$ satisfies $|\lambda|\leq\sum r_{i}$. Furthermore, using the Littlewood-Richardson rule to compute the decomposition into irreducibles (see [KT87]), one sees that the set of irreducibles appearing in the decomposition of each term (21) in degrees $k\leq 3g$ is independent of $g$; also each $S_{\langle\lambda\rangle}(V_{g})$ appearing satisfies $|\lambda|\equiv\sum r_{i}$ mod $2$. Now, Looijenga [Loo96, Theorem 1.1] has calculated the stable cohomology of $H^{*}(\mathcal{M}_{g};S_{\langle\lambda\rangle}(V_{g}))$ in degree $\leq(2g-2)/3-|\lambda|$. Since $\sum s_{i}\geq\sum r_{i}\geq|\lambda|$, we note that the contribution of each irreducible piece of each term (21) to $H^{k}(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))$ comes from Looijenga’s stable range when $k\leq(2g-2)/3$. In particular, this is independent of $g$.
@@
matching the expression shown once we identify $(V_{g}\otimes\widetilde{W}_{n})\oplus V_{g}=V_{g}\otimes W_{n}$, where $W_{n}$ is defined in (18). Finally, the stabilization argument proceeds exactly as before, except that the needed input is now $H^{*}(\mathcal{M}_{g,1},S_{\langle\lambda\rangle}(V_{g}))$, which is deduced from [Kaw08] and [RW18, Appendix B]. ∎
@@
In the cases of $\operatorname{PGL}_{2}$ and $\operatorname{SL}_{2}$, we only needed to decompose representations of the form $\wedge^{r}V_{g}$, which can be done as described in (20). However, in the case of $\operatorname{GL}_{2}$, there are also representations of the form $S_{\langle 1^{i}\rangle}(V_{g})\otimes S_{\langle 1^{j}\rangle}(V_{g})$; the relevant part of the Littlewood-Richardson rule is [KT87, p. 510]

### v2

@@ 4.1. Monodromy representations
The universal $\operatorname{PGL}_{n}$ bundle on $S_{g}\times M_{B}^{\operatorname{GL}}(S_{g,1},n,d)$ described in [21, §1] is the pullback of $\mathcal{P}$ by the quotient map $M_{B}^{\operatorname{GL}}(S_{g,1},n,d)\to\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$, as is the universal $\operatorname{PGL}_{n}$-bundle on $S_{g}\times M_{B}^{\operatorname{SL}}(S_{g,1},n,d)$ in [19, §4.1].
@@
The generation statement is a consequence of Markman’s tautological generation result for $M_{\mathrm{Dol}}^{\operatorname{GL}}(C,n,d)$ [32, Theorem 7], as reinterpreted on the Betti side by Hausel and Thaddeus [21, (5.1)]. Picking a homogeneous basis of $H^{*}(S_{g};\mathbb{Q})$, there are no relations in degrees $\leq 2(g-1)(n-1)+2$ among the images of these elements in $V_{*,j}$ by [19, Lemma 4.1.12]. ∎
@@
In later work [33, Theorem 3], Markman strengthened the generation result from rational to integral cohomology. In particular, in degrees $\leq 2(g-1)(n-1)+2$ one concludes that $H^{*}(M^{\operatorname{GL}}_{B}(S_{g,1},n,d);\mathbb{Z})$ is torsion-free. Then using the quotient description of Proposition 2.5, we conclude that the cohomology groups of the fibers $H^{*}(M^{\operatorname{SL}}_{B}(S_{g,1},n,d);\mathbb{Z}[n^{-1}])$ and $H^{*}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Z}[n^{-1}])$ are torsion-free in the same degree range. However, in order to deduce a result about integral cohomology of the universal spaces, one would also need torsion-freeness results for the various cohomology groups of $\mathcal{M}_{g}$ and $\mathcal{M}_{g,1}$ with symplectic coefficients.
@@
We follow the strategy of [21, (5.1)]. For any element of $\operatorname{Mod}_{g}$, we may choose a representative diffeomorphism $f$ of $S_{g}$ fixing a chosen point $p$. It is then straightforward to check that $\mathcal{P}$ and $f^{*}\mathcal{P}$ are both $\operatorname{PGL}_{n}$-equivariant flat $\operatorname{PGL}_{n}$-bundles over $S_{g}\times X_{g,n,d}$ trivialized over $\{p\}\times X_{g,n,d}$ such that the monodromy representation of the restriction to $S_{g}\times\{\phi\}$ is given by $\phi$, and thus $\mathcal{P}\cong f^{*}\mathcal{P}$.
@@
are canonical, and in particular invariant under homeomorphisms of $S_{g}$. (See [47] for more discussion.)
@@ 4.2. Stable rational cohomology of $\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)$
with $s_{i}=(1+2i)r_{i}^{\prime}$ for some $r_{i}^{\prime}\geq r_{i}$ satisfying $r_{i}^{\prime}\equiv r_{i}$ mod $2$. Such a term has highest weight $\sum\langle 1^{r_{i}}\rangle$, and so each irreducible piece $S_{\langle\lambda\rangle}(V_{g})$ satisfies $|\lambda|\leq\sum r_{i}$. Furthermore, using the Littlewood-Richardson rule to compute the decomposition into irreducibles (see [28]), one sees that the set of irreducibles appearing in the decomposition of each term (21) in degrees $k\leq 3g$ is independent of $g$; also each $S_{\langle\lambda\rangle}(V_{g})$ appearing satisfies $|\lambda|\equiv\sum r_{i}$ mod $2$. Now, Looijenga [30, Theorem 1.1] has calculated the stable cohomology of $H^{*}(\mathcal{M}_{g};S_{\langle\lambda\rangle}(V_{g}))$ in degree $\leq(2g-2)/3-|\lambda|$. Since $\sum s_{i}\geq\sum r_{i}\geq|\lambda|$, we note that the contribution of each irreducible piece of each term (21) to $H^{k}(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))$ comes from Looijenga’s stable range when $k\leq(2g-2)/3$. In particular, this is independent of $g$.
@@
matching the expression shown once we identify $(V_{g}\otimes\widetilde{W}_{n})\oplus V_{g}=V_{g}\otimes W_{n}$, where $W_{n}$ is defined in (18). Finally, the stabilization argument proceeds exactly as before, except that the needed input is now $H^{*}(\mathcal{M}_{g,1},S_{\langle\lambda\rangle}(V_{g}))$, which is deduced from [26] and [44, Appendix B]. ∎
@@
In the cases of $\operatorname{PGL}_{2}$ and $\operatorname{SL}_{2}$, we only needed to decompose representations of the form $\wedge^{r}V_{g}$, which can be done as described in (20). However, in the case of $\operatorname{GL}_{2}$, there are also representations of the form $S_{\langle 1^{i}\rangle}(V_{g})\otimes S_{\langle 1^{j}\rangle}(V_{g})$; the relevant part of the Littlewood-Richardson rule is [28, p. 510]